2. If the demand functions for the two consumers are:-

For consumer 1 Qd1= 10-4P1

For consumer 2 Qd2= 4-2P2

When the two prices are 3 birr.

a. Determine the market demand using the same price for the two functions?

3. The demand and supply for a 


1
Expert's answer
2022-03-09T10:31:10-0500

a. The market demand using the same price for the two functions are:

For consumer 1 Qd1= 10 - 4×3 = -2 units.

For consumer 2 Qd2 = 4 - 2×3 = -2 units.